In 20x6, Supra Ltd.’s net income computed using the variable costing method was $17007, and its net income computed using the absorption costing method was $24644. The company’s unit product cost was $7.15 under variable costing and $13.91 under absorption costing. The beginning inventory consisted of 3452 units. What was the company’s ending inventory in units? Select one:

a. 47980 units

b. 4582 units

c. 1000 units

d. 3452 units
Answer #1

Fixed manufacturing overhead cost per unit = 13.91-7.15 = 6.76

Difference in net income = 24644-17007 = 7637

Difference in inventory unit = 7637/6.76 = 1130

Ending inventory = 3452+1130 = 4582

So answer is b) 4582